### v1

@@ 3 Preliminaries
A rooted tree is a tree $T$ where one node has been chosen as the root. We remark that the root may be a leaf, and this turns out to be technically convenient for us and we will extensively use rooted trees where the root is a leaf in this paper.
@@ 8.1 Rotating hyperedges to the root
We say that a hyperedge $e\in X$ is rotatable if $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$ and there are no hyperedges $e^{\prime}$ so that $\mathcal{L}^{-1}(e^{\prime})$ is a non-child descendant of the parent $p$ of $\mathcal{L}^{-1}(e)$.
@@
  • $V_{\mathcal{T}}(\mathcal{S})=\{p,g\}$, where $p$ is the parent of $\mathcal{L}^{-1}(e)$ and $g$ is the parent of $p$.
@@
Denote the parent of $\mathcal{L}^{-1}(e)$ by $p$ and the parent of $p$ by $g$. Note that $p$ and $g$ indeed exist and are internal nodes because $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$. We construct a superbranch decomposition $\mathcal{T}^{\prime}=(T^{\prime},\mathcal{L}^{\prime})$ from $\mathcal{T}$ by contracting the edge $pg$ with the contraction basic rotation, which can be done in $2^{\mathcal{O}(k)}$ time. Denote by $t\in V(T^{\prime})$ the node of $T^{\prime}$ corresponding to $pg$. We have that $\mathcal{T}^{\prime}$ is downwards well-linked, and all nodes of $\mathcal{T}^{\prime}$ except $t$ have at most $12^{2k}+$ 1 children, while $t$ has at most $12\cdot 2^{2k}+$ 1 children.
@@
Then, we bound the increase in the potential $\Phi(\mathcal{T})$.
@@
    $\Phi(\mathcal{T}^{\prime})$ $=\Phi(\mathcal{T})+\Phi_{\mathcal{T}^{\prime}}(t)-\Phi_{\mathcal{T}}(g)-\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+(\Delta_{T}(p)+\Delta_{T}(g)-2)\cdot\log(|\mathcal{L}[g]|)-\Phi_{\mathcal{T}}(g)-\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+(\Delta_{T}(p)-1)\cdot\log(|\mathcal{L}[g]|)+(\Delta_{T}(g)-1)\cdot\log(|\mathcal{L}[g])-\Phi_{\mathcal{T}}(g)-\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+(\Delta_{T}(p)-1)\cdot\log(|\mathcal{L}[g]|)-\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+(\Delta_{T}(p)-1)\cdot\log(|\mathcal{L}[g]|)-(\Delta_{T}(p)-1)\cdot\log(|\mathcal{L}[p]|)$
    $\leq\Phi(\mathcal{T})+2^{\mathcal{O}(k)}\cdot(\log(|\mathcal{L}[g]|)-\log(|\mathcal{L}[p]|)).$
Then, we observe that
$\phi_{\mathcal{T}}(e)-\phi_{\mathcal{T}^{\prime}}(e)$ $=0pt_{T}(p)+1-\log(|\mathcal{L}[p]|)-(0pt_{T}(p)-\log(|\mathcal{L}[g]|))$
@@
Let $X^{\prime}\subseteq X$ denote the set of hyperedges $e^{\prime}\in X$ so that $\mathcal{L}^{\prime-1}(e^{\prime})$ is a child of $t$ in $T^{\prime}$. Note that the fact that $e$ is rotatable implies that $X^{\prime}$ contains all hyperedges $e^{\prime}\in X$ so that $\mathcal{L}^{\prime-1}(e^{\prime})$ is a descendant of $t$ in $T^{\prime}$.
@@
Claims 8.5 and 8.6 complete the proof that the resulting superbranch decomposition $\mathcal{T}^{\prime\prime}$ satisfies the required properties. The sequence $\mathcal{S}$ is constructed by prepending the basic rotation contracting $pg$ to the sequence $\mathcal{S}^{*}$. Clearly, $\|\mathcal{S}\|\leq 2^{\mathcal{O}(k)}+\|\mathcal{S}^{*}\|\leq 2^{\mathcal{O}(k)}$. Furthermore, we can see that $V_{\mathcal{T}}(\mathcal{S})=\{p,g\}$. ∎
@@
To argue that the total running time of the process is $2^{\mathcal{O}(k)}\log\|G\|$, it suffices to argue that in each iteration we can find a rotatable hyperedge in $X$, if one exists, efficiently. For a given hyperedge $e\in X$, we can in $\mathcal{O}(1)$ time check if $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$ by using the pointers stored in the representation of $\mathcal{T}$. We can also in $\mathcal{O}(1)$ time find the quantity $|\mathcal{L}[p]|$, where $p$ is the parent of $\mathcal{L}^{-1}(e)$. We observe that the hyperedge in $X$ that minimizes $|\mathcal{L}[p]|$ is rotatable, if any hyperedge in $X$ is rotatable, so we simply pick such a hyperedge. This runs in total time $\mathcal{O}(1)$ as $|X|\leq 3$.
Finally, we must argue that $|\mathsf{trace}_{\mathcal{T}}(\mathcal{S})|\leq 2^{\mathcal{O}(k)}\log\|G\|$, which then also implies $\|\mathcal{S}\|_{\mathcal{T}}\leq 2^{\mathcal{O}(k)}\log\|G\|$. For this, we use the following fact. If $\mathcal{S}_{i}$ is the sequence of basic rotations corresponding to the $i$-th application of Lemma 8.2, then $V_{\mathcal{T}_{i}}(\mathcal{S}_{i})\subseteq\mathsf{anc}_{T_{i}}(\mathcal{L}_{i}^{-1}(e_{i}))$, where $\mathcal{T}_{i}=(T_{i},\mathcal{L}_{i})$ is the superbranch decomposition before the $i$-th application and $e_{i}$ is the rotatable hyperedge used for that application. This implies that $\mathsf{anc}_{T_{i+1}}(\mathcal{L}_{i+1}^{-1}(e_{i}))\cap V(T)\subseteq\mathsf{anc}_{T_{i}}(\mathcal{L}_{i}^{-1}(e_{i}))\cap V(T)$, i.e., the ancestors of $e_{i}$ in the updated decomposition $\mathcal{T}_{i+1}$ that are also nodes of the initial decomposition $\mathcal{T}$ are a subset of those that are ancestors of $e_{i}$ in $\mathcal{T}_{i}$. Therefore, $\mathsf{trace}_{\mathcal{T}}(\mathcal{S})\subseteq\bigcup_{e\in X}\mathsf{anc}_{T}(\mathcal{L}^{-1}(e))$, which by the fact that $\mathcal{T}$ is $k$-good and Lemma 6.2 implies that $|\mathsf{trace}_{\mathcal{T}}(\mathcal{S})|\leq 2^{\mathcal{O}(k)}\log\|G\|$. ∎
@@ 11 Conclusions
Another direction would be to not have a pre-set treewidth bound $k$ at all, but instead let the running time of the data structure depend on the current treewidth $\mathsf{tw}(G)$. We believe that this could essentially be done, but phrasing it technically could be awkward because of the amortization. In particular, we believe that the $2^{\mathcal{O}(k)}$ factors in the running time could be replaced by $2^{\mathcal{O}(\mathsf{tw}(G))}$ factors that are not amortized, but always the current treewidth, while the $\log n$ factors would still be amortized.

### v2

@@ 3 Preliminaries
A rooted tree is a tree $T$ where one node has been chosen as the root. We remark that the root may be a leaf, and this turns out to be technically convenient for us and we will extensively use rooted trees where the root is a leaf in this paper. The parent of a non-root node $t$ in a rooted tree is the unique neighbor of $t$ on the unique path from $t$ to the root. The grandparent of $t$ (if one exists) is the parent of the parent.
@@ 8.1 Rotating hyperedges to the root
We say that a hyperedge $e\in X$ is rotatable if
  1. 1.
    $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$, and
  2. 2.
    there is no $e^{\prime}\in X$ so that $\mathcal{L}^{-1}(e^{\prime})$ is a descendant of the grandparent $g$ of $\mathcal{L}^{-1}(e)$ and $0pt_{T}(\mathcal{L}^{-1}(e^{\prime}))>0pt_{T}(\mathcal{L}^{-1}(e))$.
We then give a subroutine for rotating a rotatable hyperedge in $X$ towards the root.
@@
  • $V_{\mathcal{T}}(\mathcal{S})\subseteq\mathsf{anc}_{T}(\mathcal{L}^{-1}(X))$.
@@
Denote the grandparent of $\mathcal{L}^{-1}(e)$ by $g$, and note that it exists and is an internal node because $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$. Denote also by $X^{\prime}\subseteq X$ the set of hyperedges $e^{\prime}\in X$ so that $\mathcal{L}^{-1}(e^{\prime})$ is a descendant of $g$. The fact that $e$ is rotatable implies that for all $e^{\prime}\in X^{\prime}$, $g$ is either the parent or the grandparent of $\mathcal{L}^{-1}(e^{\prime})$. Denote by $X^{\prime\prime}\subseteq X^{\prime}$ the hyperedges $e^{\prime}\in X^{\prime}$ so that $g$ is the grandparent of $\mathcal{L}^{-1}(e^{\prime})$, and denote the set the parents of such leaves $\mathcal{L}^{-1}(e^{\prime})$ by $P^{\prime\prime}$. In particular, each $p\in P^{\prime\prime}$ is a child of $g$.
We construct a superbranch decomposition $\mathcal{T}^{\prime}=(T^{\prime},\mathcal{L}^{\prime})$ from $\mathcal{T}$ by successively contracting each edge $pg$, where $p\in P^{\prime\prime}$, with the contraction basic rotation. As $|P^{\prime\prime}|\leq|X|\leq 3$, this can be done in $2^{\mathcal{O}(k)}$ time. Let us denote the resulting node of $T^{\prime}$ by $t$. We have that $\mathcal{T}^{\prime}$ is downwards well-linked, and all nodes of $T^{\prime}$ except $t$ have at most $12^{2k}+$ 1 children, while $t$ has at most $14\cdot 2^{2k}+$ 1 children. Furthermore, $t$ is the parent of each leaf $\mathcal{L}^{\prime-1}(e^{\prime})$ so that $e^{\prime}\in X^{\prime}$, and there are no $e^{\prime}\in X\setminus X^{\prime}$ so that $\mathcal{L}^{\prime-1}(e^{\prime})$ is a descendant of $t$.
@@
Then, we bound the increase in the potential $\Phi$.
@@
    $\Phi(\mathcal{T}^{\prime})$ $=\Phi(\mathcal{T})+\Phi_{\mathcal{T}^{\prime}}(t)-\Phi_{\mathcal{T}}(g)-\sum_{p\in P^{\prime\prime}}\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+(\Delta_{T}(g)-1+\sum_{p\in P^{\prime\prime}}(\Delta_{T}(p)-1))\cdot\log(|\mathcal{L}[g]|)-\Phi_{\mathcal{T}}(g)-\sum_{p\in P^{\prime\prime}}\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+\log(|\mathcal{L}[g]|)\cdot\sum_{p\in P^{\prime\prime}}(\Delta_{T}(p)-1)-\sum_{p\in P^{\prime\prime}}\Phi_{\mathcal{T}}(p)$
    $=\Phi(\mathcal{T})+\sum_{p\in P^{\prime\prime}}(\Delta_{T}(p)-1)\cdot(\log(|\mathcal{L}[g]|)-\log(|\mathcal{L}[p]|))$
    $\leq\Phi(\mathcal{T})+2^{\mathcal{O}(k)}\cdot\sum_{p\in P^{\prime\prime}}(\log(|\mathcal{L}[g]|)-\log(|\mathcal{L}[p]|)).$
Then, we observe that when $p$ is the parent of $\mathcal{L}^{-1}(e^{\prime})$, where $e^{\prime}\in X^{\prime\prime}$,
$\phi_{\mathcal{T}}(e^{\prime})-\phi_{\mathcal{T}^{\prime}}(e^{\prime})$ $=0pt_{T}(p)+1-\log(|\mathcal{L}[p]|)-(0pt_{T}(p)-\log(|\mathcal{L}[g]|))$
@@
implying
$\phi_{\mathcal{T}}(X)-\phi_{\mathcal{T}^{\prime}}(X)\geq\sum_{p\in P^{\prime\prime}}(\log(|\mathcal{L}[g]|)-\log(|\mathcal{L}[p]|)),$
@@
Recall that $X^{\prime}\subseteq X$ is now the set of hyperedges $e^{\prime}\in X$ so that $\mathcal{L}^{\prime-1}(e^{\prime})$ is a child of $t$ in $T^{\prime}$, and $X^{\prime}$ contains all hyperedges $e^{\prime}\in X$ so that $\mathcal{L}^{\prime-1}(e^{\prime})$ is a descendant of $t$ in $T^{\prime}$.
@@
Claims 8.5 and 8.6 complete the proof that the resulting superbranch decomposition $\mathcal{T}^{\prime\prime}$ satisfies the required properties. The sequence $\mathcal{S}$ is constructed by prepending the basic rotation contracting $pg$ to the sequence $\mathcal{S}^{*}$. Clearly, $\|\mathcal{S}\|\leq 2^{\mathcal{O}(k)}+\|\mathcal{S}^{*}\|\leq 2^{\mathcal{O}(k)}$. Furthermore, we can see that $V_{\mathcal{T}}(\mathcal{S})\subseteq\mathsf{anc}_{T}(\mathcal{L}^{-1}(X))$. ∎
@@
To argue that the total running time of the process is $2^{\mathcal{O}(k)}\log\|G\|$, it suffices to argue that in each iteration we can find a rotatable hyperedge in $X$, if one exists, efficiently. For a given hyperedge $e\in X$, we can in $\mathcal{O}(1)$ time check if $0pt_{T}(\mathcal{L}^{-1}(e))\geq 3$ by using the pointers stored in the representation of $\mathcal{T}$. We can also in $\mathcal{O}(1)$ time find the quantity $|\mathcal{L}[g]|$, where $g$ is the grandparent of $\mathcal{L}^{-1}(e)$. We observe that the hyperedge in $X$ that minimizes $|\mathcal{L}[g]|$ is rotatable, if any hyperedge in $X$ is rotatable, so we simply pick such a hyperedge. This runs in total time $\mathcal{O}(1)$ as $|X|\leq 3$.
Finally, we must argue that $|\mathsf{trace}_{\mathcal{T}}(\mathcal{S})|\leq 2^{\mathcal{O}(k)}\log\|G\|$, which then also implies $\|\mathcal{S}\|_{\mathcal{T}}\leq 2^{\mathcal{O}(k)}\log\|G\|$. For this, we use the following fact. If $\mathcal{S}_{i}$ is the sequence of basic rotations corresponding to the $i$-th application of Lemma 8.2, then $V_{\mathcal{T}}(\mathcal{S}_{i})\subseteq\mathsf{anc}_{T_{i}}(\mathcal{L}_{i}^{-1}(X))$, where $\mathcal{T}_{i}=(T_{i},\mathcal{L}_{i})$ is the superbranch decomposition before the $i$-th application This implies that $\mathsf{anc}_{T_{i+1}}(\mathcal{L}_{i+1}^{-1}(X))\cap V(T)\subseteq\mathsf{anc}_{T_{i}}(\mathcal{L}_{i}^{-1}(X))\cap V(T)$, i.e., the ancestors of $\mathcal{L}_{i+1}^{-1}(X)$ in the updated decomposition $\mathcal{T}_{i+1}$ that are also nodes of the initial decomposition $\mathcal{T}$ are a subset of those that are ancestors of $\mathcal{L}_{i}^{-1}(X)$ in $\mathcal{T}_{i}$. Therefore, $\mathsf{trace}_{\mathcal{T}}(\mathcal{S})\subseteq\mathsf{anc}_{T}(\mathcal{L}^{-1}(X))$, which by the fact that $\mathcal{T}$ is $k$-good and Lemma 6.2 implies that $|\mathsf{trace}_{\mathcal{T}}(\mathcal{S})|\leq 2^{\mathcal{O}(k)}\log\|G\|$. ∎
@@ 11 Conclusions
Another direction would be to not have a pre-set treewidth bound $k$ at all, but instead let the running time of the data structure depend on the current treewidth $\mathsf{tw}(G)$. We believe that with minor modifications, our data structure already achieves something along these lines, but phrasing it formally would get technical because of the amortization.